Given the total cost function TC = 18q2

+ 4q + 16,200

(a) Find marginal cost (MC) and average coat (AC) as functions of q

(b) Show that when MC < AC, AC is falling, and when MC > AC, AC is rising


1
Expert's answer
2021-04-06T07:08:28-0400

The marginal cost function can be found as follows:



"MC=\dfrac{dTC}{dQ}=36q+4."


The average cost can be found as follows:



"AC=\dfrac{TC}{q}=18q+4+\dfrac{16200}{q}."


If MC < AC, then AC is falling until MC = AC. Let's equate MC and AC:



"36q+4=18q+4+\dfrac{16200}{q},""18q^2-16200=0,""q=30."

If "q>30", MC>AC, therefore, AC is rising.